# Does provable absence of barren plateaus imply classical simulability? Or, why we need to rethink variational quantum computing

### v1

@@ Abstract
A large amount of effort has recently been put into understanding the barren plateau phenomenon. In this perspective article, we face the increasingly loud elephant in the room and ask a question that has been hinted at by many but not explicitly addressed: Can the structure that allows one to avoid barren plateaus also be leveraged to efficiently simulate the loss classically? We present strong evidence that commonly used models with provable absence of barren plateaus are also classically simulable, provided that one can collect some classical data from quantum devices during an initial data acquisition phase. This follows from the observation that barren plateaus result from a curse of dimensionality, and that current approaches for solving them end up encoding the problem into some small, classically simulable, subspaces. This sheds serious doubt on the non-classicality of the information processing capabilities of parametrized quantum circuits for barren plateau-free landscapes and on the possibility of superpolynomial advantages from running them on quantum hardware. We end by discussing caveats in our arguments, the role of smart initializations, and by highlighting new opportunities that our perspective raises.
@@ I Introduction
In recent years, the initial excitement attracted by variational quantum algorithms [1, 2, 3] and quantum machine learning [4, 5, 6, 7, 8, 9] has been tempered by the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Namely, there is a growing awareness that a large class of quantum learning architectures exhibit loss function landscapes that concentrate exponentially in system size towards their mean value. On such landscapes, exponential resources are required for training, prohibiting the successful scaling of variational quantum algorithms. Hence, identifying architectures and training strategies that provably do not lead to barren plateaus has become a highly active area of research. Examples of such strategies include shallow circuits with local measurements [22, 13, 23, 24, 25], dynamics with small Lie algebras [20, 47, 48, 49, 50, 51, 55], identity initializations [36, 52, 57], embedding symmetries into the circuit’s architecture [58, 59, 60, 61, 62, 63], and certain classes of quantum generative models [30, 29, 26].
@@
Our general argument is supported by an analysis of widely used schemes through which we show that all considered methods for avoiding barren plateaus can be efficiently classically simulated. Fundamentally, it is the very proof of absence of barren plateaus that allows us to identify the polynomially-sized subspaces in which the relevant part of the computation lives. Using this information, we can then determine the set of expectation values one needs to estimate (either classically or quantumly) to enable classical simulations.
@@ II Definitions for barren plateaus and simulability
Over the past few years, there has been a tremendous amount of work put forward to understand if the loss functions in a given problem class are trainable. Several sources of untrainability have been detected [7] such as the presence of sub-optimal local minima [69, 70, 71, 72, 20] and expressivity limitations [73]. However, the vast majority of trainability analysis has been concentrated around the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 29, 26, 27, 28, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. When a problem exhibits a barren plateau, its loss function becomes—on average—exponentially concentrated with the system size [74, 22]. We say a class $\mathcal{C}$ of problems is provably barren plateau-free, i.e., $\mathcal{C}\in\overline{{\rm BP}}$, if one can show that
@@
First, let us define the task of computing a loss function, such as that in Eq. (1), for a given problem. We will say that an algorithm can compute the loss functions in an instance $\mathcal{I}$ if, with high probability for $\bm{\theta}\sim\mathcal{P}$, it can implement a function $\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$, i.e.,
$\left(\tilde{\ell}_{\bm{\theta}}(\rho,O)-\ell_{\bm{\theta}}(\rho,O)\right)^{2}\leqslant\epsilon\,.$ (3)
@@
As previously mentioned, the above definition guarantees that we can compute the loss function with high probability when the parameters are sampled according to $\mathcal{P}$. One might also be interested in computing the loss given any parameter settings. Thus, we will also consider a stronger version where an algorithm can compute the loss function in an instance $\mathcal{I}$ if for all $\bm{\theta}$, it can implement a function $\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$.
@@
We also consider classical simulation algorithms enhanced by polynomial-size data obtained from quantum experiments, which we denote as Classical Simulation enhanced with Quantum Experiments (${\rm CSIM}_{\mathrm{QE}}$). In this case, one is given a problem instance and is allowed to use a quantum computer for an initial data acquisition phase, which takes no more than polynomial time. During this phase, one can prepare copies of the initial quantum state, apply some operations, and independently measure them via some efficient tomographic or classical shadow techniques [77, 65, 78]. Such a procedure can be used to obtain an efficiently storable classical representation (i.e., storable in string of size in $\mathcal{O}(\operatorname{poly}(n))$) of the state, unitary or measurement operator. Once this initial phase is over, one cannot access the quantum device anymore, and the computation of $\ell_{\bm{\theta}}(\rho,O)$ has to be done purely classically. A problem $\mathcal{C}$ is in ${\rm CSIM}_{\mathrm{QE}}$ if a polynomial-time classical algorithm, which can utilize data obtained from quantum devices in an initial data acquisition phase, can compute every instance in $\mathcal{C}$ [^1] Note that the problem of estimating a loss with data from a quantum computer can be cast into a decision problem. In this case, $\rm CSIM$} can be related to BPP, ${\rm CSIM}_{\mathrm{QE}}$ is closely connected to BPP/Samp [64] and even more closely connected to BPP/qgenpoly [79], and $\rm QSIM$} is related to BQP. .
@@ III What leads to absence of barren plateaus?
$S_{n}$-equivariant [63] Permutation equivariant (P) $O\in\mathcal{B}_{\lambda}$, Tr[$\Tr[\rho_{\lambda}^{2}]\in\Omega(1/\operatorname{poly}(n))$
Matchgate circuit [51] Product of $\mathcal{O}(1)$ Majoranas (P) $O\in\mathcal{B}_{\lambda}$, Tr[$\Tr[\rho_{\lambda}^{2}]\in\Omega(1/\operatorname{poly}(n))$
@@
Small Lie algebra $\mathfrak{g}$ [49, 61, 55] Operators in $\mathfrak{g}$ (P) $O\in i\mathfrak{g}$, Tr[$\Tr[\rho_{\mathfrak{g}}^{2}]\in\Omega(1/\operatorname{poly}(n))$
@@
Provably barren plateau-free architectures live in classically identifiable polynomial subspaces.
@@ IV Connection between absence of barren plateaus and simulability
To wrap up, we highlight that combining Claim 1 and Claim 2 with our case-by-case analysis we can arrive at the conclusion: In none of the standard problem instances with provable absence of barren plateaus does the parametrized quantum circuit need to be implemented on a quantum computer in order to estimate the loss in polynomial time. Note that our results do not necessarily imply a dequantization of variational quantum computing as a whole, since a quantum device might still be needed for the initial data acquisition phase. Still, in a way, Claims 1 and 2 do dequantize the variational part of the model and shed some serious doubts on the non-classicality of the information processing being done by a barren plateau-free loss function.
@@ V.2 New opportunities
As discussed in the previous section, we know that if the problem has effective subspaces, then one could potentially simulate and train the loss given some initial set of parameters, but that the loss might not be faithful for the full optimization procedure (see Fig. 5(a)). To mitigate this issue, we envision a novel form of hybrid quantum computation as shown in Fig. 5(b). Here, there is still a feedback loop between the classical and quantum device (similar to that of standard variational methods), but instead of updating parameters in a quantum circuit, one updates the data acquired from the initial state. As such, as the optimization progresses, one uses the quantum computer to take new measurements on the initial state, and then uses this information for a more faithful, updated classical simulation. While we have not tested the validity of this protocol, we believe it might be useful and encourage the community to try it out.
@@ A.1 Small dynamical Lie algebras
Next, one needs to obtain the projection of $\rho$ and $O$ into the dynamical Lie algebra (denoted as $\rho_{\mathfrak{g}}$ and $O_{\mathfrak{g}}$, respectively), which requires computing the inner product of the initial state and the measurement operator with the elements of a basis of $i\mathfrak{g}$. Such computation will generally be case dependent, but or all cases considered we have always found an efficient tomographic procedure (see below for more information). From here, we can employ a recently introduced classical simulation algorithm called $\mathfrak{g}$-sim [125, 93], which allows us to (back-)propagate the elements of the dynamical Lie algebra with unitaries of the Lie group. For this algorithm to work, one needs to compute the structure constants of the algebra. Importantly, computing these constants takes a time that scales polynomially in $\dim(\mathfrak{g})$. However, this suffices to show that
@@
Here we find it important to note that we have assumed above that the Lie algebra is simple. However, $\mathfrak{g}$ will generally decompose as a direct sum of simple Lie algebras (plus an abelian component) as $\mathfrak{g}=\oplus_{\eta}\mathfrak{g}_{\eta}$. Here, barren plateaus can be avoided (and thus simulations are enabled) if one of the simple components is polynomially large and, if $\rho$ and $O$ have large projections therein.
@@ A.1.1 $U(1)$-equivariant circuit
Several proofs of absence of barren plateaus for $U(1)$-equivariant have been presented in the literature [20, 47, 48]. Crucially, they all somewhat rely on the fact that the restriction of the dynamical Lie algebra into the subspace with $k$ Hamming weight can be, at most, of dimension $\mathcal{O}(\binom{n}{k}^{2})$, meaning that a necessary condition for the loss function to not exhibit barren plateaus is that the initial state’s Hamming weight is not too large. Note that bitstrings with Hamming weight nkn-k can trivially be mapped to bitstrings of Hamming weight $k$ by flipping all bits, and, while our following statements equally apply to Hamming weight nkn-k, we will focus on the case of Hamming weight $k\leqslant n/2$. As such, one finds
@@ A.4.1 Hamiltonian variational ansatz
$\mathcal{C}_{{\rm HVA}}\subset\overline{{\rm BP}}\,,\,\,\text{if $\rho$ follows an area law of entanglement}.$ (34)
@@ A.4.2 Gaussian initialization in deep variational quantum circuits
$\mathcal{C}_{{\rm GI}}\subset\overline{{\rm BP}}\,,\,\,\text{if $\rho$ follows an area law of entanglement}\,.$ (37)
@@ A.5.1 Standard QCNN with tracing-out pooling layers
$\mathcal{C}_{{\rm QCNN}}\subset\overline{{\rm BP}}\,,\,\,\text{if $\rho$ follows an area law of entanglement}.$ (40)
@@ A.5.2 Precision-limited QCNN with controlled measurements pooling layers
copies of $\rho$. Here, $\left\lVert\cdot\right\rVert$ is the operator norm. By Ref. [77], this classical shadows description of $\rho$ is of size $\operatorname{poly}\left(N\right)$. Theorem S1 combined with Proposition S1 of [77] guarantees that, with probability at least $1-\delta$, this classical shadows description suffices to estimate the expectation values of
@@
[…]
@@ Appendix B Examples of non-concentrated but also non-simulable loss functions
[…]
@@
However, this example is rather contrived. For one, it considers an initialization over a discrete set of parameters. This is a long way from the spirit of more variational quantum algorithms. Moreover, this discrete initialization is doing much of the heavy lifting in this example. To see this one can also consider a variant of the above problem where the parameters are sampled uniformly in the range $0,2\pi]$. We will call this distribution $\mathcal{P}_{\rm cont.}$ and the corresponding problem class $\mathcal{C}_{{\rm Bool}}^{\rm cont.}$.

### v2

@@ Abstract
A large amount of effort has recently been put into understanding the barren plateau phenomenon. In this perspective article, we face the increasingly loud elephant in the room and ask a question that has been hinted at by many but not explicitly addressed: Can the structure that allows one to avoid barren plateaus also be leveraged to efficiently simulate the loss classically? We present strong evidence that commonly used models with provable absence of barren plateaus are also classically simulable, provided that one can collect some classical data from quantum devices during an initial data acquisition phase. This follows from the observation that barren plateaus result from a curse of dimensionality, and that current approaches for solving them end up encoding the problem into some small, classically simulable, subspaces. Thus, while stressing quantum computers can be essential for collecting data, our analysis sheds serious doubt on the non-classicality of the information processing capabilities of parametrized quantum circuits for barren plateau-free landscapes. We end by discussing caveats in our arguments, the role of smart initializations and the possibility of provably superpolynomial, or simply practical, advantages from running parametrized quantum circuits.
@@ I Introduction
In recent years, the initial excitement attracted by variational quantum algorithms [1, 2, 3] and quantum machine learning [4, 5, 6, 7, 8, 9] has been tempered by the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Namely, there is a growing awareness that a large class of quantum learning architectures exhibit loss function landscapes that concentrate exponentially in system size towards their mean value. On such landscapes, exponential resources are required for training, prohibiting the successful scaling of variational quantum algorithms. Hence, identifying architectures and training strategies that provably do not lead to barren plateaus has become a highly active area of research. Examples of such strategies include shallow circuits with local measurements [22, 13, 23, 24, 25, 26], dynamics with small Lie algebras [20, 48, 49, 50, 51, 52, 56], identity initializations [37, 53, 58], embedding symmetries into the circuit’s architecture [59, 60, 61, 62, 63, 64], and certain classes of quantum generative models [31, 30, 27].
@@
Our general argument is supported by an analysis of widely used schemes through which we show that all considered methods for avoiding barren plateaus can be efficiently classically simulated. Fundamentally, it is the very proof of absence of barren plateaus that allows us to identify the polynomially-sized subspaces in which the relevant part of the computation lives. Using this information, we can then determine the set of expectation values one needs to estimate (either classically or quantumly) to enable classical simulations. In the latter case the quantum computer is still essential, but is accessed non-adaptively to generate a classical surrogate rather than via a hybrid-quantum classical optimization loop.
@@ II Definitions for barren plateaus and simulability
Over the past few years, there has been a tremendous amount of work put forward to understand if the loss functions in a given problem class are trainable. Several sources of untrainability have been detected [7] such as the presence of sub-optimal local minima [70, 71, 72, 73, 20] and expressivity limitations [74]. However, the vast majority of trainability analysis has been concentrated around the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 30, 27, 28, 29, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. When a problem exhibits a barren plateau, its loss function becomes—on average—exponentially concentrated with the system size [75, 22]. We say a class $\mathcal{C}$ of problems is provably barren plateau-free, i.e., $\mathcal{C}\in\overline{{\rm BP}}$, if one can show that
@@
First, let us define the task of computing a loss function, such as that in Eq. (1), for a given problem. We will say that an algorithm can compute the loss functions in an instance $\mathcal{I}$ if, with high probability for $\bm{\theta}\sim\mathcal{P}$, it can implement a function ~$\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$, i.e.,
(~$\left(\tilde{\ell}_{\bm{\theta}}(\rho,O)-\ell_{\bm{\theta}}(\rho,O)\right)^{2}\leqslant\epsilon\,.$ (3)
@@
As previously mentioned, the above definition guarantees that we can compute the loss function with high probability when the parameters are sampled according to $\mathcal{P}$. One might also be interested in computing the loss given any parameter settings. Thus, we will also consider a stronger version where an algorithm can compute the loss function in an instance $\mathcal{I}$ if for all $\bm{\theta}$, it can implement a function ~$\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$.
@@
We also consider classical simulation algorithms enhanced by polynomial-size data obtained from quantum experiments, which we denote as Classical Simulation enhanced with Quantum Experiments (${\rm CSIM}_{\mathrm{QE}}$). In this case, one is given a problem instance and is allowed to use a quantum computer for an initial data acquisition phase, which takes no more than polynomial time. During this phase, one can prepare copies of the initial quantum state, apply some operations, and independently measure them via some efficient tomographic or classical shadow techniques [78, 66, 79]. Such a procedure can be used to obtain an efficiently storable classical representation (i.e., storable in string of size in $\mathcal{O}(\operatorname{poly}(n))$) of the state, unitary or measurement operator. Once this initial phase is over, one cannot access the quantum device anymore, and the computation of $\ell_{\bm{\theta}}(\rho,O)$ has to be done purely classically. A problem $\mathcal{C}$ is in ${\rm CSIM}_{\mathrm{QE}}$ if a polynomial-time classical algorithm, which can utilize data obtained from quantum devices in an initial data acquisition phase, can compute every instance in $\mathcal{C}$ [^1] Note that the problem of estimating a loss with data from a quantum computer can be cast into a decision problem. In this case, ${\rm CSIM}$ can be related to BPP, ${\rm CSIM}_{\mathrm{QE}}$ is closely connected to BPP/Samp [65] and even more closely connected to BPP/qgenpoly [80], and ${\rm QSIM}$ is related to BQP. Speaking more informally, algorithms in ${\rm CSIM}_{\mathrm{QE}}$ are also often described as classical surrogate methods..
@@ III What leads to absence of barren plateaus?
[…]
@@
[…]
@@
Standard provably barren plateau-free architectures live in classically identifiable polynomial subspaces.
@@ IV Connection between absence of barren plateaus and simulability
To wrap up, we highlight that combining Claim 1 and Claim 2 with our case-by-case analysis we can arrive at the conclusion: In none of the standard problem instances with provable absence of barren plateaus does the parametrized quantum circuit need to be implemented on a quantum computer in order to estimate the loss in polynomial time. Note that our results do not imply a dequantization of variational quantum computing as a whole, since a quantum device might still be needed for the initial data acquisition phase. Viewed this way, our results can be viewed positively as highlighting the potential of a different learning paradigm where the quantum computer is used non-adaptively to create a classical surrogate of the loss landscape. Still, in a way, Claims 1 and 2 do dequantize the variational part of the model and shed some serious doubts on the non-classicality of the information processing being done by a barren plateau-free loss function.
@@ V.2 New opportunities
As discussed in the previous section, we know that if the problem has effective subspaces, then one could potentially simulate and train the loss given some initial set of parameters, but that the loss might not be faithful for the full optimization procedure (see Fig. 5(a)). To mitigate this issue, we envision a novel form of hybrid quantum computation as shown in Fig. 5(b). Here, there is still a feedback loop between the classical and quantum device (similar to that of standard variational methods), but instead of updating parameters in a quantum circuit, one updates the data acquired from the initial state. As such, as the optimization progresses, one uses the quantum computer to take new measurements on the initial state, and then uses this information for a more faithful, updated classical simulation. While we have not tested the performance of this protocol, we believe it might be useful and encourage the community to try it out.
@@ A.1 Small dynamical Lie algebras
Next, one needs to obtain the projection of $\rho$ and $O$ into the dynamical Lie algebra (denoted as $\rho_{\mathfrak{g}}$ and $O_{\mathfrak{g}}$, respectively), which requires computing the inner product of the initial state and the measurement operator with the elements of a basis of $i\mathfrak{g}$. Such computation will generally be case dependent, but or all cases considered we have always found an efficient tomographic procedure (see below for more information). From here, we can employ a recently introduced classical simulation algorithm called $\mathfrak{g}$-sim [126, 94], which allows us to (back-)propagate the elements of the dynamical Lie algebra with unitaries of the Lie group. For this algorithm to work, one needs to compute the structure constants of the algebra. Importantly, computing these constants takes a time that scales polynomially in $\dim(\mathfrak{g})$. Thus, this suffices to show that
@@
Here we find it important to note that we have assumed above that the Lie algebra is simple. However, $\mathfrak{g}$ will generally decompose as a direct sum of simple Lie algebras (plus an abelian component) as $\mathfrak{g}=\oplus_{\eta}\mathfrak{g}_{\eta}$. Here, barren plateaus can be avoided (and thus simulations are enabled) if at least one of the simple components is polynomially large, and $\rho$ and $O$ have large projections therein.
@@ A.1.1 $U(1)$-equivariant circuit
Several proofs of absence of barren plateaus for $U(1)$-equivariant have been presented in the literature [20, 48, 49]. Crucially, they all somewhat rely on the fact that the restriction of the dynamical Lie algebra into the subspace with $k$ Hamming weight can be, at most, of dimension $\mathcal{O}(\binom{n}{k}^{2})$, meaning that a necessary condition for the loss function to not exhibit barren plateaus is that the initial state’s Hamming weight is not too large. Note that bitstrings with Hamming weight $n-k$ can trivially be mapped to bitstrings of Hamming weight $k$ by flipping all bits, and, while our following statements equally apply to Hamming weight $n-k$, we will focus on the case of Hamming weight $k\leqslant n/2$. As such, one finds
@@ A.4.1 Hamiltonian variational ansatz
[…]
@@ A.4.2 Gaussian initialization in deep variational quantum circuits
[…]
@@ A.5.1 Standard QCNN with tracing-out pooling layers
[…]
@@ A.5.2 Precision-limited QCNN with controlled measurements pooling layers
[…]
@@
$O_{m_{1},\ldots,m_{n-k}}=O\otimes\ket{m_{1}}\bra{m_{1}}\otimes\ldots\otimes\ket{m_{n-k}}\otimes\bra{m_{n-k}}.$ | . (48)
@@ Appendix B Examples of non-concentrated but also non-simulable loss functions
$\ell_{\bm{\theta}}(O)=\bra{0^{n}}\left(\bigotimes_{i=1}^{n}e^{i\theta_{i}X}\right)U^{\dagger}Z_{1}U\left(\bigotimes_{i=1}^{n}e^{-i\theta_{i}X}\right)\ket{0^{n}}\,,$ (52)
where $O=Z_{1}$ and we have explicitly taken the initial state $\rho$ to be the all-zero state. Let us further suppose that each $\theta_{i}$ is uniformly sampled from the discrete set $\{0,\pi/2,\pi,3\pi/2,2\pi\}$ such that the parametrized quantum circuit prepares a random computational basis state. That is, $\left(\bigotimes_{i=1}^{n}e^{-i\theta_{i}X}\right)\ket{0^{n}}=\ket{x}$ where $|x\rangle$ is a random bitstring. We call this distribution $\mathcal{P}_{\rm disc}$. We will then define $\mathcal{C}_{{\rm Bool}}^{\rm disc.}$ as the class of all instances where the circuit is of the form in Eq. (50) where $U$ implements a classically hard Boolean function (even with advice strings) and $\bm{\theta}\sim\mathcal{P}_{\rm disc}$.
@@
However, this example is rather contrived. For one, it considers an initialization over a discrete set of parameters. This is a long way from the spirit of more variational quantum algorithms. Moreover, this discrete initialization is doing much of the heavy lifting in this example. To see this one can also consider a variant of the above problem where the parameters are sampled uniformly in the range $[0,2\pi]$. We will call this distribution $\mathcal{P}_{\rm cont.}$ and the corresponding problem class $\mathcal{C}_{{\rm Bool}}^{\rm cont.}$.